x/3 - y/2 = - 10/3
I Need Help With This Elimination Method Problem Involving Fractions. Fractions Confuse me. Fractions are Not My friend!!!!!!!!!!!!!

Answer by fractalier(6550)

From x/3 - y/2 = - 10/3, we can clear all of the fractions in one step by multiplying all of the terms by the lowest common denominator...in this case, 6.
So we have
6(x/3) - 6(y/2) = 6(-10/3)
2x - 3y = -20
If you are looking to graph this, solve it for y and get
y = (2/3)x + 20/3